OpenStudy (albert0898):

if x^2-y^2=12 and x-y=4 what is the value of x^2+2xy+y^2 a. 3 b. 8 c. 9 d. 12 e. 16

myininaya (myininaya):

you could factor x^2-y^2 and then use what x-y is equal to to find x+y

myininaya (myininaya):

then recall x^2+2xy+y^2 can be written as (x+y)^2

OpenStudy (albert0898):

Erm.... Well here's what I did so far... I did (x-y)^2 and got x^2 - 2xy + y^2 = 16 I don't know what else to do

myininaya (myininaya):

well can you try what I said?

myininaya (myininaya):

x^2-y^2 can be factored

myininaya (myininaya):

it is a difference of squares

myininaya (myininaya):

you are given the value of x-y this will allow you to find x+y

myininaya (myininaya):

and then all you have to do after finding x+y is find (x+y)^2

myininaya (myininaya):

So have you factored x^2-y^2 yet?

OpenStudy (albert0898):

(x-y)(x+y)

myininaya (myininaya):

\[x^2-y^2=12 \\ (x-y)(x+y)=12 \\ \text{ you are given } x-y=4 \\ \text{ just plug in } \\ 4(x+y)=12\] \[\text{ and you want to find } (x+y)^2\]

OpenStudy (albert0898):

Ahh!! Now I understand it. Thank you!

myininaya (myininaya):

k great